Find the frequency of a tuning fork that takes 5.15 ms to complete one oscillation.

Round your answer to 2 decimal places.

Pls do not forget the unit


1
Expert's answer
2022-05-17T08:57:54-0400
"f=\frac{1}{T}=\frac{1}{5.15*10^{-3}\rm s}=194.17\:\rm Hz"